# Decoding Speculative Decoding

### v3

@@ Abstract
Speculative Decoding is a widely used technique to speed up inference for Large Language Models (LLMs) without sacrificing quality. When performing inference, speculative decoding uses a smaller draft model to generate speculative tokens and then uses the target LLM to verify those draft tokens. The speedup provided by speculative decoding heavily depends on the choice of the draft model. In this work, we perform a detailed study comprising over 350 experiments with LLaMa-65B and OPT-66B using speculative decoding and delineate the factors that affect the performance gain provided by speculative decoding. Our experiments indicate that the performance of speculative decoding depends heavily on the latency of the draft model, and the draft model’s capability in language modeling does not correlate strongly with its performance in speculative decoding. Based on these insights we explore a new design space for draft models and design hardware-efficient draft models for speculative decoding. Our newly designed draft model for LLaMa-65B can provide 111% higher throughput than existing draft models and can generalize further to the LLaMa-2 model family and supervised fine-tuned models.
@@
Minghao Yan1thanks: Correspondence: Minghao Yan <myan@cs.wisc.edu>, Saurabh Agarwal1, Shivaram Venkataraman1 1University of Wisconsin-Madison
Figure 1: This figure shows the speculative decoding process. In vanilla LLM inference, after the prompt is processed into KV caches (Prefill phase), LLM generates the output token by token in an autoregressive manner (Autoregressive generation phase). In speculative decoding, a draft model is first used to generate $n$ candidate tokens at each step (Draft token generation phase). The LLM verifies the candidate tokens and accepts $k$ ($k\leq n$) tokens (LLM verification phase). Since LLM knows all $n$ candidate tokens in advance, this step is identical to a prefill step of length $n$. In both cases, this process is repeated until either an end-of-sequence (EOS) token is generated or the maximum generation limit has been reached.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as a cornerstone of modern computational linguistics, offering unprecedented capabilities in generating and interpreting human language. As the demand for faster and more efficient language processing grows, understanding and optimizing the inference throughput of these models becomes increasingly crucial. Decoder-only LLMs Brown et al. (2020); Touvron et al. (2023a, b) use auto-regressive decoding to perform inference. Auto-regressive decoding is known to be hardware inefficient Miao et al. (2023); Liu et al. (2023a), leading to poor resource utilization and low throughput during inference.
@@
Several methods Yu et al. (2022); Wang et al. (2020); Kwon et al. (2023); Dao et al. (2023); Hong et al. (2023) have been studied to optimize the serving of LLMs. One promising approach to improve the throughput for serving LLMs without accuracy loss is speculative decoding Stern et al. (2018); Xia et al. (2023a); Leviathan et al. (2023). When using speculative decoding to serve an LLM (usually 100s of billion parameters), a draft model (a significantly smaller LLM) is used to generate speculative tokens. The target LLM model then verifies the output of the draft model and only outputs tokens that match its output. In the case of speculative decoding, the target LLM for inference acts as a verifier for the draft model. By leveraging faster inference of smaller draft models, speculative decoding turns autoregressive decoding on the target LLM into a more hardware-friendly batched operation (similar to “prefill”), thereby increasing throughput while preserving accuracy.
@@
Given the promised benefits of speculative decoding, this paper first focuses on understanding the key factors that dictate the throughput improvements that can be obtained. We perform a comprehensive benchmarking study and profile speculative decoding to characterize bottlenecks. We perform over 350 experiments, using LLMs like LLaMa-65B, OPT-66B, and fine-tuned chat models such as Vicuna-33B Chiang et al. (2023) as target models and LLaMa and OPT families as draft models, ranging from 5$\approx$ 5$\times$ 5 × to 528×528$\times$528 × fewer parameters than the target models. Our findings show that the key bottleneck in speculative decoding is the draft model’s latency, highlighting the need to optimize draft model designs.
@@
Next, we find that existing draft models, which are typically designed only for improving accuracy in a given parameter budget, are sub-optimal for maximizing the throughput with speculative decoding. From our experiments, we first observe that the draft model latency is bottlenecked by model depth, and higher model depth leads to increased latency (Section 3.2). Second, we also find that draft model accuracy on language modeling tasks does not correlate strongly with its performance in speculative decoding (Section 3.3), i.e., a draft model with higher accuracy on language modeling task can have similar TAR to a model with lower accuracy. Based on these two insights, we propose designing new draft models that trade increased depth for width (thus retaining the same parameter count) and show that our new draft models can boost inference throughput using speculative decoding by over 60%. Finally, we show how pruning methods like Sheared-LLaMa Xia et al. (2023b) can generate smaller draft models with favorable configurations.
We have open-sourced our code  [^1] https://github.com/uw-mad-dash/decoding-speculative-decoding and distilled models [^2] https://huggingface.co/minghaoyan/ on Github and HuggingFace.
@@
  • We show the need for a systematic redesign of draft models used for speculative decoding. We demonstrate that using accuracy on language modeling tasks to choose the draft model for speculative decoding can lead to suboptimal choices, and our experiments highlight that redesigning draft models can improve the throughput of speculative decoding by up to 111% in sampling-based decoding and 60% in greedy decoding. Based on these insights, our pruned LLaMa-796M provides up to $111\%$ higher throughput than Sheared-LLaMa-1.3B while using only $0.8\%$ of tokens (0.42B vs 50.42B) used to train Sheared-LLaMa-1.3B. We also show that LLaMa-796M works well for other LLMs, such as LLaMa-2 families of models and supervised fine-tuned models (Vicuna-33B).
  • Finally, we also study how improvements in models and hardware can further impact draft model design for future generations of LLMs. (Section 5.1).
@@ 2.1 Background
A decoder-only LLM performs inference in two phases: a prefill phase and an autoregressive-decoding phase. In the prefill phase, the LLM is initialized with a context or prompt, formulated as $C=\{c_{1},c_{2},...,c_{n}\}$, where $C$ represents the input context and $n$ the length of the prefill. In the prefill phase, the model processes the whole input context in parallel and performs next-word prediction. During the autoregressive-decoding phase, the model generates new text sequentially, one token at a time, building upon the context provided in the prefill phase. Due to its sequential nature, the autoregressive decoding phase is widely known to be memory bandwidth bound on modern GPUs Leviathan et al. (2023).
To improve hardware utilization and throughput, Schuster et al. (2022); Chen et al. (2023) proposed speculative decoding, where a significantly smaller draft model generates multiple tokens, and the target LLM performs verification on the generated tokens in parallel. The verification is akin to prefill stage in LLM inference. As long as more than one token is accepted on average, speculative decoding can potentially provide speedups. Figure 1 shows how inference using speculative decoding differs from auto-regressive decoding. It is widely reported Miao et al. (2023); Liu et al. (2023a) that the number of tokens accepted by the target model influences the speedup provided by speculative decoding.
@@ 2.2 Related Work
LLM Inference
There has been significant amount of work on improving LLM serving including work in Orca Yu et al. (2022), LightSeq Wang et al. (2020), DeepSpeed Inference Aminabadi et al. (2022), PagedAttention Kwon et al. (2023), FlashDecoding Dao et al. (2023) and FlashDecoding++ Hong et al. (2023). These works seek to improve LLM inference by better utilization of hardware. There are lines of work that have looked at pruning LLMs based on input context to speed up inference Liu et al. (2023b) or using shallower and wider neural networks for machine translation Kasai et al. (2020). However, in this work, we focus on speculative decoding Leviathan et al. (2023); Chen et al. (2023); Santilli et al. (2023), which has been inspired by speculative execution in hardware Hennessy and Patterson (2011).
Speculative Decoding
Several prior works have studied ways to improve speculative decoding. Liu et al. (2023a) seeks to continuously train the draft model on the output of the target model to improve the token acceptance rate. However, training on the same hardware during inference can be challenging, depending on the inference request rate and hardware utilization. Predictive Pipeline Decoding (PPD) Yang et al. (2023) was one of the first methods to introduce the use of early exit Schuster et al. (2022) from the target model to obtain draft tokens. Similar to PPD, Draft&Verify Zhang et al. (2023) seeks to combine the use of early exit with speculative decoding, where the early exit Schuster et al. (2022); Bae et al. (2023) from the target model acts as a draft token. A drawback of these methods is that the maximum benefit in latency is capped. For example, in speculative decoding, we can use draft models that are orders of magnitude (e.g., $\approx$100x-1000x) smaller than the target model, while early exit methods usually exit after performing inference over at least a fourth of the model Schuster et al. (2022), thus, limiting the gain in throughput. Other lines of work, such as Medusa Cai et al. (2024), propose fine-tuning multiple generation heads within the LLM that do not match the LLM output distribution exactly but maintain the generation quality.
@@
Figure 3: Performance Breakdown of OPT speculative decoding, lookahead length is set to be optimal for each draft model found empirically.
@@ 3 Understanding Speculative Decoding
To study the effects of the choice of the draft model, we first perform a detailed study on serving OPT-65B and LLaMa-65B (two popular LLMs) using speculative decoding.
Setup.
We implement speculative decoding in the Microsoft Deepspeed library Microsoft (2023). We use the same setup as SpecInfer Miao et al. (2023), first using the draft model to generate draft tokens and then using the target model to verify the output of the draft model. We set the batchsize to 1 and use greedy decoding. For all our experiments, we use 4 Nvidia 80GB A100 GPUs. We perform our experiment on the OPT and LLaMa base models Zhang et al. (2022); Touvron et al. (2023a) on MMLU Hendrycks et al. (2020), Hellaswag Zellers et al. (2019), and Chatbot Arena datasets Zheng et al. (2023). For MMLU, we use the standard 5-shot setup. The remaining datasets were evaluated in a zero-shot setting. Note that since our goal is to test our draft model’s ability to study the target model’s behavior, we do not instruct the model to emit a single-letter answer to MMLU and Hellaswag questions, but instead opt for an open-ended generation approach where the model can provide as much as explanation as it sees fit, which aligns better with a real-world chatbot setting. For MMLU, we feed the model with the question and choices; for Hellaswag, we feed the model with incomplete sentences without the choices. We use OPT-66B and LLaMa-65B as the target LLM for OPT and LLaMa series and use OPT-125M, OPT-350M, OPT-1.3B, OPT-2.7B, and OPT-6.7B as draft models for OPT series, and LLaMa-7B, and LLaMa-13B as draft models for LLaMa series.
Metrics.
To quantify the performance of different draft models when performing inference on a target model, we measure throughput (tokens generated per second) and TAR (Figure 2). We note that the primary goal of speculative decoding is to improve throughput.
3.1 Bottlenecks in Speculative Decoding
In Figure 3, the time taken by the draft model for token generation increases with an increase in model sizes, going from 6.23 ms for OPT-125M to 18.56 ms for OPT-6.7B. However, even the smallest draft model, OPT-125M, still takes significant time in a speculative decoding iteration to perform draft model autoregressive decoding. Though the target LLM has a higher latency in each decoding iteration, it only has to perform one prefill operation on the entire candidate token sequence. In contrast, the draft model has to perform multi-step autoregressive decoding sequentially, creating a bottleneck. This highlights why draft model latency is one of the key bottlenecks in speculative decoding performance. We note that while Figure 3 uses lookahead values from 6 to 8, depending on the draft model, even if we scale lookahead length to hundreds of tokens, the target model verification time stays constant. The draft model latency remains the bottleneck due to the difference in efficiency between prefill and auto-regressive decoding. Next, we investigate how we can reduce draft model latency.
@@ 3.2 Understanding Draft Model Latency
Figure 4 shows the results of these three benchmarks. In the first benchmark (Figure 4(a)), we vary the number of attention heads, feed-forward dimension, and layers in a model to keep the model parameters at around 350M. The detailed configuration for each model can be found in Table 7 in the Appendix. The plot shows that the autoregressive decoding latency is linear in terms of layer depth despite each model having roughly the same parameter count.
@@ 3.3 Understanding Draft Model TAR
In prior work Leviathan et al. (2023), speculative decoding throughput is modeled by $\frac{1-\alpha^{\gamma+1}}{(1-\alpha)(\gamma c+1)}$, where $\frac{1-\alpha^{\gamma+1}}{1-\alpha}$ represents the improvement factor (expected number of tokens matched in each iteration) and $\gamma c+1$ represents the combined latency of draft and target models. Therefore, tokens matched per iteration (also known as TAR) have a linear effect on speculative decoding throughput.
In this section, we perform experiments to understand the correlation between the accuracy of a model on popular NLP tasks and its TAR. We plot the accuracy of a model against the TAR it achieves in Figure 5. Surprisingly, we find that TAR correlates little to the model’s accuracy on a task. We believe this lack of correlation is due to the majority of tokens in a sentence not being content words Chen et al. (2023), which do not affect the model’s accuracy on a specific task.
For example, if a user asks a model: What is the capital of Uruguay? An LLM may correctly answer: The capital of Uruguay is Montevideo. But a draft model, without retaining this much knowledge, may respond incorrectly: The capital of Uruguay is Paris. For model accuracy evaluation, this would be a failure. However, this would be a good set of candidate tokens in speculative decoding, as the first five words are generated correctly. Therefore, as shown in Figure 5, TAR increases sub-linearly with an increase in model size, irrespective of its accuracy on the task. Results on more datasets can be found in the Appendix (Figure 10).
@@
(a) LLaMa Series on MMLU
(b) LLaMa Series on Hellaswag

### v4

@@ Abstract
Speculative Decoding is a widely used technique to speed up inference for Large Language Models (LLMs) without sacrificing quality. When performing inference, speculative decoding uses a smaller draft model to generate speculative tokens and then uses the target LLM to verify those draft tokens. The speedup provided by speculative decoding heavily depends on the choice of the draft model. In this work, we perform a detailed study comprising over 350 experiments with LLaMa-65B and OPT-66B using speculative decoding and delineate the factors that affect the performance gain provided by speculative decoding. Our experiments indicate that the performance of speculative decoding depends heavily on the latency of the draft model, and the draft model’s capability in language modeling does not correlate strongly with its performance in speculative decoding. Based on these insights we explore a new design space for draft models and design hardware-efficient draft models for speculative decoding. Our newly designed draft model can provide 111% higher throughput than existing draft models and our approach generalizes further to all LLaMa models (1/2/3.1) and supervised fine-tuned models.
@@
Minghao Yanthanks: Correspondence: Minghao Yan <myan@cs.wisc.edu>,  Saurabh Agarwal,  Shivaram Venkataraman Department of Computer Sciences University of Wisconsin-Madison
Figure 1: This figure shows the speculative decoding process. In vanilla LLM inference, after the prompt is processed into KV caches (Prefill), LLM generates the output token by token (Autoregressive generation). In speculative decoding, a draft model is first used to generate $n$ candidate tokens at each step (Draft token generation). The LLM verifies the candidate tokens and accepts $k$ ($k\leq n$) tokens (LLM verification).
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as a cornerstone of modern computational linguistics, offering unprecedented capabilities in generating and interpreting human language. As the demand for faster and more efficient language processing grows, understanding and optimizing the inference throughput of these models becomes increasingly crucial. Decoder-only LLMs Brown et al. (2020); Touvron et al. (2023a, b) use autoregressive decoding to perform inference. Autoregressive decoding is known to be hardware inefficient Miao et al. (2023); Liu et al. (2023a), leading to poor resource utilization and low throughput during inference.
@@
Several methods Yu et al. (2022); Wang et al. (2020); Kwon et al. (2023); Dao et al. (2023); Hong et al. (2023) have been studied to optimize the serving of LLMs. One promising approach to improve the throughput for serving LLMs without accuracy loss is speculative decoding Stern et al. (2018); Xia et al. (2023a); Leviathan et al. (2023). When using speculative decoding to serve an LLM (usually 10s to 100s of billion parameters), a draft model (a significantly smaller LLM) is used to generate speculative tokens. The target LLM model then verifies the output of the draft model and only outputs tokens that match its output. In the case of speculative decoding, the target LLM for inference acts as a verifier for the draft model. By leveraging faster inference of smaller draft models, speculative decoding turns autoregressive decoding on the target LLM into a more hardware-friendly batched operation (similar to “prefill”), thereby increasing throughput while preserving accuracy.
@@
Given the promised benefits of speculative decoding, this paper first focuses on understanding the key factors that dictate the throughput improvements that can be obtained. We perform a comprehensive benchmarking study and profile speculative decoding to characterize bottlenecks. We perform over 350 experiments, using LLMs like LLaMa-65B, OPT-66B, and fine-tuned chat models such as Vicuna-33B Chiang et al. (2023) as target models and LLaMa and OPT families as draft models, ranging from 5$\approx$ 5$\times$ 5 × to 528×528$\times$528 × fewer parameters than the target models. Our findings show that the key bottleneck in speculative decoding is the draft model’s latency (time to generate candidate tokens from the draft model), highlighting the need to study the design of draft models for speculative decoding.
@@
While studying the design of draft models, we observed two interesting phenomena: First we observe that the draft model latency is bottlenecked by model depth, and higher model depth leads to increased latency (Section 3.2) even with the same number of parameters. Secondly, we observe that draft model accuracy on language modeling tasks does not correlate strongly with its performance in speculative decoding (Section 3.3), i.e., a draft model with higher accuracy on language modeling task can have similar TAR to a model with lower accuracy. These two phenomena show that existing draft models used for speculative decoding were primarily designed only to achieve maximum accuracy for a given parameter budget and are sub-optimal for maximizing the throughput with speculative decoding.
Based on these two insights, we propose designing new draft models that trade increased depth for width (thus retaining the same parameter count) and show that our new draft models can boost inference throughput using speculative decoding by over 60%. Finally, we show how pruning methods like Sheared-LLaMa Xia et al. (2023b) can be used to generate smaller draft models with favorable configurations on three different families of models OPT, LLaMa, and LLaMa-3.1.
@@
  • We show a systematic redesign of draft models used for speculative decoding is needed to maximize the efficiency of speculative decoding. We demonstrate that using accuracy on language modeling tasks to choose the draft model for speculative decoding leads to suboptimal choices. By redesigning draft models, we improved speculative decoding throughput by up to 111%.
  • We show that our design leads to a $37\%$ reduction in KV-Caches, enabling larger batch sizes, and also outperforms other popular methods like self-speculative decoding Zhang et al. (2023) in several different setups.
@@ 2.1 Background
A decoder-only LLM performs inference in two phases: prefill and autoregressive decoding. In the prefill phase, the LLM is initialized with a context or prompt, formulated as $C=\{c_{1},c_{2},...,c_{n}\}$, where $C$ represents the input context and $n$ the length of the prefill. In the prefill phase, the model processes the whole input context in parallel and performs next-word prediction. During the autoregressive decoding, the model generates new text sequentially, a token at a time, building upon the context provided in the prefill phase. Due to its sequential nature, the autoregressive decoding phase is widely known to be memory bandwidth bound on modern GPUs Leviathan et al. (2023).
To improve hardware utilization and throughput,  Leviathan et al. (2023) and Chen et al. (2023) proposed speculative decoding, where a smaller draft model generates multiple tokens, and the target LLM verifies the generated tokens in parallel. The verification is akin to prefill stage in LLM inference. As long as more than one token is accepted on average, speculative decoding can potentially provide speedups. Figure 1 shows how inference using speculative decoding differs from autoregressive decoding. It is widely reported Miao et al. (2023); Liu et al. (2023a) that the number of tokens accepted by the target model affects the speedup provided by speculative decoding.
@@ 2.2 Related Work
LLM Inference: There has been significant amount of work on improving LLM serving including work in Orca Yu et al. (2022), LightSeq Wang et al. (2020), DeepSpeed Inference Aminabadi et al. (2022), PagedAttention Kwon et al. (2023), FlashDecoding Dao et al. (2023) and FlashDecoding++ Hong et al. (2023). These works seek to improve LLM inference by better utilization of hardware. Other lines of work have looked at pruning LLMs based on input context to speed up inference Liu et al. (2023b) or using shallower and wider neural networks for machine translation Kasai et al. (2020). However, in this work, we focus on speculative decoding Leviathan et al. (2023); Chen et al. (2023); Santilli et al. (2023), which has been inspired by speculative execution in hardware Hennessy and Patterson (2011).
Speculative Decoding: Several prior works have studied ways to improve speculative decoding.  Liu et al. (2023a) seeks to continuously train the draft model on the output of the target model to improve the token acceptance rate. However, training on the same hardware during inference can be challenging. Predictive Pipeline Decoding (PPD) Yang et al. (2023) introduced the use of early exit Schuster et al. (2022) from the target model to obtain draft tokens. Similar to PPD, Draft&Verify Zhang et al. (2023) seeks to combine the use of early exit with speculative decoding, where the early exit Schuster et al. (2022); Bae et al. (2023) from the target model acts as a draft token. A drawback of these methods is that the maximum benefit in latency is capped. For example, in speculative decoding, we can use draft models that are orders of magnitude (e.g., $\approx$100x-1000x) smaller than the target model, while early exit methods usually exit after performing inference over at least a fourth of the model Schuster et al. (2022), thus, limiting the gain in throughput. Other lines of work, such as Medusa Cai et al. (2024), propose fine-tuning multiple generation heads within the LLM that do not match the LLM output distribution exactly but maintain the generation quality. In addition, other works have looked at improving speculative decoding via learning an encoder-decoder-based draft model Xia et al. (2023a), generating multiple draft tokens Sun et al. (2024b); Yang et al. (2024), draft model distillation Zhou et al. (2024), or focusing on long-context scenarios Sun et al. (2024a); Chen et al. (2024).
@@ 3 Understanding Speculative Decoding
To study the effects of the choice of the draft model, we first perform a detailed study on serving OPT-65B and LLaMa-65B using speculative decoding.
Setup: We implement speculative decoding in the Microsoft Deepspeed library Microsoft (2023). We use the same setup as SpecInfer Miao et al. (2023), first using the draft model to generate draft tokens and then using the target model to verify the output of the draft model. We set the batch size to 1 and use greedy decoding. For all our experiments, we use 4 Nvidia 80GB A100 GPUs. We perform our experiment on the OPT and LLaMa base models Zhang et al. (2022); Touvron et al. (2023a) on MMLU Hendrycks et al. (2020), Hellaswag Zellers et al. (2019), and Chatbot Arena datasets Zheng et al. (2023). For MMLU, we use the standard 5-shot setup. The remaining datasets were evaluated in a zero-shot setting. Note that since our goal is to test our draft model’s ability to study the target model’s behavior, we do not instruct the model to emit a single-letter answer to MMLU and Hellaswag questions, but instead opt for an open-ended generation approach where the model can provide as much as explanation as it sees fit, which aligns better with a real-world chatbot setting. For MMLU, we feed the model with the question and choices; for Hellaswag, we feed the model with incomplete sentences without the choices. We use OPT-66B and LLaMa-65B as the target LLM for OPT and LLaMa series and use OPT-125M, 350M, 1.3B, 2.7B, and 6.7B variants as draft models for OPT series, and LLaMa-7B and 13B as draft models for LLaMa series.
Metrics: To quantify the performance of different draft models when performing inference on a target model, we measure throughput (tokens generated per second) and TAR (Figure 2). We note that the primary goal of speculative decoding is to improve throughput.
Figure 4: This figure shows the performance breakdown of speculative decoding on OPT models, look ahead length is set to be optimal for each draft model found empirically.
@@
3.1 Bottlenecks in Speculative Decoding
In Figure 4, the time taken by the draft model for each token generation step increases with an increase in model sizes, going from 6.23 ms for OPT-125M to 18.56 ms for OPT-6.7B (Table 11 in the Appendix). However, even the smallest draft model, OPT-125M, still takes significant time in a speculative decoding iteration to perform draft model autoregressive decoding. Though the target LLM has a higher latency in each decoding iteration, it only has to perform one prefill operation on the entire candidate token sequence. In contrast, the draft model has to perform multi-step autoregressive decoding sequentially, creating a bottleneck. This highlights why draft model latency is one of the key bottlenecks in speculative decoding performance. We note that while Figure 4 uses look ahead values (the number of tokens generated by the draft model) from 6 to 8, depending on the draft model, even if we scale look ahead length to hundreds of tokens, the target model verification time stays constant. The draft model latency remains the bottleneck due to the difference in efficiency between prefill and autoregressive decoding. Next, we investigate how to reduce draft model latency.
@@ 3.2 Understanding Draft Model Latency
Figure 3 shows the results of these three benchmarks. In the first benchmark (Figure 3(a)), we vary the number of attention heads, feed-forward dimension, and layers in a model to keep the model parameters at around 350M. The detailed configuration for each model can be found in Table 13 in the Appendix. The plot shows that autoregressive decoding latency scales linearly with layer depth despite similar total model parameters.
@@ 3.3 Understanding Draft Model TAR
In prior work Leviathan et al. (2023), speculative decoding throughput is modeled by $\frac{1-\alpha^{\gamma+1}}{(1-\alpha)(\gamma c+1)}$, where $\frac{1-\alpha^{\gamma+1}}{1-\alpha}$ represents the improvement factor (expected number of tokens matched in each iteration) and $\gamma c+1$ represents the combined latency of draft and target models. Therefore, tokens accepted per iteration (also known as TAR) have a linear effect on speculative decoding throughput.
In this section, we perform experiments to understand the correlation between the accuracy of a model on popular NLP tasks and its TAR. We plot the accuracy of a model against the TAR it achieves in Figure 5. Surprisingly, we find that TAR correlates little to the model’s accuracy on a task. We believe this lack of correlation is due to the majority of tokens in a sentence not being content words Chen et al. (2023), which do not affect the model’s accuracy on a specific task. Results on more datasets can be found in the Appendix (Figure 8).